OpenStudy (anonymous):

Solve 3|2x + 2| − 7 = 5 for x.

OpenStudy (anonymous):

1) add 7 to both sides 2) divide both sides by 3 what do you get?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yes, on the right you get 4, on the left you have \(|2x+2|\) so you have to solve \[|2x+2|=4\] which you do by solving two separate equations: \[2x+2=-4\] or \[2x+2=4\]